A man owns stock valued at $2000. One day the stock drops by 7% and then gains the same percentage back the next day. How much money is the stock valued at the end of the second day?
Found 2 solutions by mananth, MathTherapy:
Answer by mananth(16946)

A man owns stock valued at $2000. One day the stock drops by 7% and then gains the same percentage back the next day. How much money is the stock valued at the end of the second day?
A man owns stock valued at $2000.
drops by 7%
value = 2000-7%*2000
=2000-140 = 1860
On 1860 it increases by 7%
Value next day = 1860 +7% *1860
=1860+130.20 =1980.20
the stock valued at $1980.20

Answer by MathTherapy(10552)


A man owns stock valued at $2000. One day the stock drops by 7% and then gains the same percentage back the next day. How much money is the stock valued at the end of the second day?
Correct answer: 
Decrease in value: 100 - 7 = 93%, or .93
Increase in value: 100 + 7 = 107%, or 1.07